OpenStudy (ammarah):

Solve the system using the elimination method.. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ At a local market, 3 apples and 2 pears cost $2.70. Three apples cost the same as 4 pears. How much do apples and pears cost?

OpenStudy (akashdeepdeb):

Let apples cost = x let pears cost = y So one equation would be = 3x + 2y = 2.70 The other equation = 3x = 4y Now eliminate and solve. :D

OpenStudy (ammarah):

wait isnt it 3x+4y for the second equation?

OpenStudy (ammarah):

and wouldnt the second equation also be equal to 2.70?

OpenStudy (ammarah):

wait thats the second equation?

OpenStudy (ammarah):

so its 3x + 2y =2.70 4y +2y=2.70

OpenStudy (ammarah):

wait i mean 4x

OpenStudy (ammarah):

u put 4y...